# SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation

### v2

@@ 4.3 Inference Performance
Batch Inference Performance. Fig. 3 shows the results of Llama-3.1-8B and Llama-3.1-70B across several workloads with a range of input lengths. SwiftKV achieves higher combined throughput than the baseline model across all the workloads we evaluated.
For Llama-3.1-8B-Instruct, with 2K input tokens per prompt, SwiftKV achieves 1.21.3×1.2-1.3$\times$1.2 - 1.3 × higher combined throughput than the baseline model, and our benefits increase further to 1.81.9×1.8-1.9$\times$1.8 - 1.9 × higher combined throughput with 128K inputs. Note that for an input length of 8K tokens, SwiftKV achieves a staggering 30K tokens/sec/GPU (480 TFLOPS/GPU). For Llama-3.1-70B with 2K input tokens per prompt, SwiftKV achieves 1.41.5×1.4-1.5$\times$1.4 - 1.5 × higher combined throughput than the baseline model, which improves to 1.82.0×1.8-2.0$\times$1.8 - 2.0 × better combined throughput for 128K inputs. [^5] While the total compute savings is roughly 2×2$\times$2 ×, the end-to-end speedup is lower due to two main reasons: i) the performance improvement is limited to the decoding computation which needs the output activation of all the layers. Fig. 2 (right) shows the max possible speedup for Llama-3.1-8B-Instruct during model forward pass despite the decoding overhead, and ii) due to additional vLLM overheads outside of the model forward pass, such as sampling, optimizing which is beyond the scope of the paper. As expected, SwiftKV provides greater improvements when the inputs are long.
@@
Figure 3: Combined input and output throughput for Llama-3.1-8B (left) and Llama-3.1-70B (right) across a range of input lengths (bottom). For each experiment, a number of requests are sent to vLLM at once such that the total number of tokens is roughly 15M. Each request generates 256 output tokens.
@@
Interactive-Inference Performance. Fig. 4 shows the TTFT and TPOT of Llama-3.1-70B-Instruct across a range of request arrival rates and input lengths. When the arrival rate is too high, the TTFT explodes due to the request queue accumulating faster than they can be processed by the system. However, SwiftKV can sustain 1.52.0×1.5-2.0$\times$1.5 - 2.0 × higher arrival rates before experiencing such TTFT explosion. When the arrival rate is low, SwiftKV can reduce the TTFT by up to 50% for workloads with longer input lengths. In terms of TPOT, SwiftKV achieves significant reductions for all but the lowest arrival rates, up to 60% for certain settings. A similar story unfolds for Llama-3.1-8B, which can be found in Fig. B.1 in the Appendix.
5 Ablation and Discussion
5.1 Compute Reduction vs Memory Compression
A key aspect of SwiftKV is combining prefill compute reduction (SingleInputKV) and KV cache compression (AcrossKV). While many prior works address KV cache compression alone, they are only effective in scenarios with limited GPU memory, and can have limited impact on recent datacenter GPUs (e.g., A100 and H100) with sufficient memory and inference is compute-bound.
To illustrate, we consider an “ideal” KV compression scheme, where every layer’s KV cache is merged into a single layer (Merge-all-Layers). We retain the computation for all KV operations (i.e., $W_{kv}^{T}X$) but eliminate the memory for all layers > 1, leading to a single layer of KV cache. Merge-all-Layers represents a “best case compression scenario” with (1) extreme compression ratio beyond any published technique, e.g. 32×32$\times$32 × and 80×80$\times$80 × for Llama-3.1 8B and 70B, respectively, and (2) zero overhead, while most techniques (e.g., quantization, low-rank decomposition) add extra computations or data conversions.
Table 5 shows the throughput attained by Merge-all-Layers compared with the baseline model and its SwiftKV variants under various memory constraints. As shown, Merge-all-Layers outperforms only in very low memory scenarios (e.g. 16GB and 20GB) when there is barely enough memory for just the model weights, and is only marginally (10%) better than the baseline model when using all 80GB memory. On the other hand, SingleInputKV attains 35% higher throughput than the baseline at 80GB even without any AcrossKV. When combined with 4×4$\times$4 × AcrossKV using FP8-quantized KV cache, SwiftKV can approach the throughput of Merge-all-Layers even at a more limited 20GB of memory.
Throughput (tokens/s) Memory Baseline Merge-all- 50% SingleInputKV 50% SingleInputKV 50% SingleInputKV Layers + 4×4$\times$4 × AcrossKV + 4×4$\times$4 × AcrossKV (FP8) 80GB 22.9K 25.1K 31.0K 31.2K 32.0K 40GB 20.6K 25.2K 27.3K 28.4K 28.9K 20GB 10.8K 25.2K 12.2K 18.0K 23.2K 16GB OOM 24.8K OOM 4.22K 7.28K
@@ 5.1 Compute Reduction vs Memory Compression
Table 6: Impact of Distillation and Full/Partial Model Finetuning on Llama-3.1-8B-Instruct
@@
Setting Arc-Challenge Winogrande Hellaswag TruthfulQA MMLU MMLU-CoT GSM-8K Avg. 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots (a) The effect of distillation W/o Distillation 79.44 77.27 78.71 51.14 65.55 65.60 72.71 70.06 W Distillation 80.38 78.22 79.30 54.54 67.30 69.73 79.45 72.70 (b) Full model finetuning vs. part model finetuning Full Model 76.79 74.82 76.42 53.08 62.94 64.20 69.37 68.23 Part Model 80.38 78.22 79.30 54.54 67.30 69.73 79.45 72.70
@@ 5.2 The impact of distillation
To demonstrate the effectiveness of our distillation, we train Llama-3.1-8B-Instruct with 50% SingleInputKV and no AcrossKV using the standard language model loss, and compare it with our distillation based approach discussed in  Sec. 3.4. The results are shown in Table 6 (a). As we can see, the model trained with distillation has a 2.64 point higher average. Particularly, for generative tasks, i.e., MMLU-Cot and GSM-8K, the performance improvement is 4.13 and 6.74, respectively.
@@
Our distillation method only fine-tuned the ${\mathbf{W}}_{QKV}$ parameters, as discussed in Sec. 3.4, with the hypothesis that it preserves the knowledge from the original models compared to fine-tuning the entire model. This hypothesis aligns with Meng et al. (2024), Geva et al. (2021), and Elhage et al. (2021), which suggest that MLP layers player a more prominent role in storing knowledge.
To validate this, we fine-tuned a model with 50% SingleInputKV on Llama-3.1-8B-Instruct where all parameters in the latter 50% of layers are trained. The results are shown in Table 6 (b). The model quality of full model distillation is about 4.5 points lower than our proposed partial model distillation.
5.3 Combining with other KV cache compression methods
SwiftKV explores an orthogonal design space from many KV cache compression methods, which means that it can be easily combined with them, e.g., sliding window (Jiang et al., 2023), token-level pruning (Liu et al., 2024d), quantization (Hooper et al., 2024) etc. In this section, we show the combined effect of SwiftKV with per-token KV cache FP8 quantization  (Yao et al., 2022) using PyTorch’s natively supported float8_e4m3fn.  Table 7 shows the accuracy degradation is within 0.4 points for all cases, even though we applied post-training quantization with no quantization-aware finetuning.
Appendix C explores a second, potentially interesting, trade-off between AcrossKV (inter-layer) vs GQA (intra-layer) KV cache sharing.
Table 7: Llama-3.1-8B-Instruct KV cache quantization results.
5.4 Other Ablation Studies
We show additional ablations that explore the impact of the distillation dataset in Appendix D, and how SwiftKV can enable simple early-exiting for decode tokens in Appendix E.
@@ 6 Conclusions
In this paper, we presented SwiftKV, a novel model transformation for reducing inference cost for prompt-dominant workloads, combined with a KV cache reduction strategy to reduce memory footprint, and a light-weight distillation procedure to preserve model accuracy. While we presented strong results on the effectiveness of SwiftKV, exploration of parameter-preserving model transformations for inference optimization is still in its early stages. We have identified both limitations as well as areas of improvement. Given the simplicity and effectiveness of SwiftKV, we hope that this will spark further exploration which we hope will continue to lower the latency and cost of inference.
Limitations and Future Work
It is important for every work to acknowledge its limitations and suggest future directions, particularly for LLM-related works. In our work, we did not aim to optimize the training data selection though we provide potential ways in Sec. D. Additionally, we did not include a detailed benchmark analysis for our method. However, as shown in Sec. D, we ensured that our datasets were not cherry-picked to overfit the reported tasks. Furthermore, we did not finetune our model with advanced post-training approaches, like DPO and RLHF, which we leave for future work. Finally, we hypothesize that our method can work even better when combined with pretraining or continued-pretraining, but due to resources constraints, we did not explore this direction. We hope to revisit these ideas in the future.
Appendix A Additional Motivation
Figure A.1: The input similarity of small scale models (left) and large scale models (right).
Appendix B Experimental Details
B.1 Training and evaluation details
We directly use the Huggingface LLama-3.1 checkpoints, particularly, “meta-llama/Meta-Llama-3.1-8B-Instruct” and “meta-llama/Meta-Llama-3.1-70B-Instruct”. For datasets, we use the supervised finetuning datasets from “HuggingFaceH4/ultrachat_200k” and “teknium/OpenHermes-2.5”, which in total is about 1.2M samples, and about 160M tokens. We set training epochs to be 2, learning rate to be 3e-4, weight decay to be 0.05, warm up ratio to be 5%, maximum sequence length to be 8192 with attention separated sequence packing, the distillation temperature to be 2.0, and the training batch size to be 32 for both Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct.
Our evaluation follows https://huggingface.co/neuralmagic/Meta-Llama-3.1-8B-Instruct-FP8 using the github repository https://github.com/neuralmagic/lm-evaluation-harness/tree/llama_3.1_instruct. The main reason behind this is that the implemention from the repository aligns with original Llama-3.1 evaluation, which has superme scores over the original Lm-eval-harness repository. One issue we found in the provided commands is the one used to run MMLU-5-shots. Directly using the command does not give us desired accuracy. Therefore, we added both --apply_chat_template and --fewshot_as_multiturn, and the accuracy is even slightly higher than what they reported.
For all tasks, we follow the same number of few shots and/or chain of thoughts as the provided commands. We present the number of shots and metrics used in the paper in Table B.1.
Table B.1: The setting for different tasks
@@ B.1 Training and evaluation details
B.2 Inference Speedup Evaluation Details
Figure B.1: Time to first token (TTFT, top) and time per output token (TPOT, bottom) for input lengths 2000 (left), 8000 (middle), and 32000 (right) for Llama-3.1-8B. For each experiment, a range of different request arrival rates is simulated. Each request generates 256 output tokens.
@@ B.2 Inference Speedup Evaluation Details
Llama-3.1-8B Latency Evaluation.
See Fig. B.1.
B.3 Additional Model Evaluations
Full per-task evaluation scores for Llama-3.2-3B-Instruct, Llama-3.1-405B-Instruct (FP8), Mistral-Small-Instruct-2409, and Deepseek-V2-Lite-Chat can be found in Table B.2, Table B.3, Table B.4, and Table B.5, respectively.
Table B.2: Llama-3.2-3B-Instruct
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 75.17 68.59 73.32 51.45 62.01 62.48 72.32 66.47 SwiftKV ✓(25%) 75.59 69.77 72.34 52.80 61.89 62.39 71.11 66.55 SwiftKV ✓(40%) 75.34 68.98 71.37 51.10 61.80 61.62 68.68 65.55 SwiftKV ✓(50%) 71.25 68.75 70.77 51.29 59.63 59.94 67.02 64.09 SwiftKV ✓(40%) 2-way (25%) 74.82 68.66 71.41 50.67 61.55 61.03 67.77 65.13 SwiftKV ✓(40%) 4-way (37.5%) 75.59 69.21 70.79 50.89 61.35 60.82 67.70 65.19
Table B.3: Llama-3.1-405B-Instruct (FP8)
Model SingleInputKV AcrossKV Arc-Challenge Winogrande Hellaswag TruthfulQA MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 94.7 87.0 88.3 64.7 87.5 88.1 96.1 86.6 SwiftKV ✓(50%) 94.0 86.3 88.1 64.2 85.7 87.5 95.2 85.9
Table B.4: Mistral-Small-Instruct-2409
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 84.12 84.68 87.27 56.85 73.33 74.86 86.50 78.23 SwiftKV ✓(25%) 84.04 84.84 87.03 55.97 72.88 74.69 85.21 77.80 SwiftKV ✓(50%) 83.53 83.97 86.30 55.63 72.91 74.04 84.30 77.24 SwiftKV ✓(50%) 2-way (25%) 83.36 84.05 86.22 56.20 72.30 73.70 84.68 77.21 SwiftKV ✓(50%) 4-way (37.5%) 82.93 83.82 86.17 56.00 72.29 73.00 82.48 76.66
Table B.5: Deepseek-V2-Lite-Chat
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 65.53 74.66 81.56 50.98 56.86 50.61 68.69 64.12 SwiftKV ✓(25%) 65.44 75.05 81.52 50.53 56.91 50.92 68.99 64.19 SwiftKV ✓(45%) 65.61 73.95 80.82 50.20 56.33 51.56 66.11 63.51 SwiftKV ✓(45%) 2-way (25%) 65.52 74.26 80.23 49.85 55.59 50.51 65.57 63.07 SwiftKV ✓(45%) 4-way (37.5%) 61.34 75.21 79.80 48.39 54.82 30.80 64.89 59.32
Appendix C Inter-layer AcrossKV vs Intra-Layer KV cache Reduction

### v3

@@ 4.2 Inference Performance
We focus on two common production scenarios:
  1. 1.
    Batch-Inference: When processing requests in bulk or serving a model under high usage demand, it is important to achieve high combined throughput in terms of input and output tokens to cost-effectively serve the model.
  2. 2.
    Interactive-Inference: In interactive scenarios (e.g., chatbots, copilots), metrics that define the end-user experience are the time-to-first-token (TTFT) and time-per-output-token (TPOT). Low TTFT and TPOT are desirable to deliver smooth usage experiences.
We evaluate the end-to-end inference performance using Llama-3.1-8B-Instruct running on 1 NVIDIA H100 GPU with 80GB of memory, Llama-3.1-70B-Instruct running on 4 NVIDIA H100 GPUs with 4-way tensor parallelism. We show results using vLLM and refer to our SGLang results in Appendix A.4, and provide the full hardware and vLLM configurations in Appendix A.2.
Batch Inference Performance.
Fig. 3 shows the results of Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct across several workloads with a range of input lengths. SwiftKV achieves higher combined throughput than the baseline across all the workloads we evaluated. For Llama-3.1-8B-Instruct, with 2K input tokens per prompt, SwiftKV achieves 1.21.3×1.2-1.3$\times$1.2 - 1.3 × higher combined throughput than the baseline, and our benefits increase further to 1.81.9×1.8-1.9$\times$1.8 - 1.9 × higher combined throughput with 128K inputs. Note that for an input length of 8K tokens, SwiftKV achieves a staggering 30K tokens/sec/GPU (480 TFLOPS/GPU). For Llama-3.1-70B-Instruct with 2K input tokens per prompt, SwiftKV achieves 1.41.5×1.4-1.5$\times$1.4 - 1.5 × higher combined throughput than the baseline, which improves to 1.82.0×1.8-2.0$\times$1.8 - 2.0 × better combined throughput for 128K inputs.
Throughput (tokens/s) Memory Baseline Merge-all- 50% SwiftKV 50% SwiftKV 50% SwiftKV Layers + 4×4$\times$4 × AcrossKV + 4×4$\times$4 × AcrossKV (FP8) 80GB 22.9K 25.1K 31.0K 31.2K 32.0K 40GB 20.6K 25.2K 27.3K 28.4K 28.9K 20GB 10.8K 25.2K 12.2K 18.0K 23.2K 16GB OOM 24.8K OOM 4.22K 7.28K
Figure 3: Combined input and output throughput for Llama-3.1-8B-Instruct (left) and Llama-3.1-70B-Instruct (right) across input lengths (bottom). Roughly 15M tokens worth of requests are sent for each experiment, and each request generates 256 output tokens.
@@
Interactive-Inference Performance.
@@
Fig. 4 shows the TTFT and TPOT of Llama-3.1-70B-Instruct across a range of request arrival rates and input lengths, and we refer to Fig. A.1 in the Appendix for Llama-3.1-8B-Instruct. When the arrival rate is too high, the TTFT explodes due to the request queue accumulating faster than they can be processed by the system. However, SwiftKV can sustain 1.52.0×1.5-2.0$\times$1.5 - 2.0 × higher arrival rates before experiencing such TTFT explosion. When the arrival rate is low, SwiftKV can reduce the TTFT by up to 50% for workloads with longer input lengths. In terms of TPOT, SwiftKV achieves significant reductions for all but the lowest arrival rates, up to 60% for certain settings.
At first, it may be counter-intuitive that SwiftKV can reduce TPOT by only optimizing the prefill compute and not decode compute. However, in most open-source inference systems today, including vLLM and SGLang, prefill and decode are run on the same GPUs, whether they be interleaved Yu et al. (2022) or mixed Holmes et al. (2024); Agrawal et al. (2024). This means prefill and decode may contend for GPU time, and reducing prefill compute also benefits decode latency.
Inference on Real-World Requests.
In Appendix A.5, we evaluate SwiftKV on SGLang using real-world requests from ShareGPT ShareGPT Team (2023), which are collected in the wild from users of ChatGPT OpenAI (2022). We show that the throughput improvements due to SwiftKV transfer well to real-world length distributions.
5 Ablations and Discussions
@@
Table 4: Impact of Distillation and Full/Partial Model Finetuning on Llama-3.1-8B-Instruct.
@@
Setting Arc-Challenge Winogrande Hellaswag TruthfulQA MMLU MMLU-CoT GSM-8K Avg. 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots (a) The effect of distillation W/o Distill 79.44 77.27 78.71 51.14 65.55 65.60 72.71 70.06 W Distill 80.38 78.22 79.30 54.54 67.30 69.73 79.45 72.70 (b) Full model finetuning vs. part model finetuning Full Model 76.79 74.82 76.42 53.08 62.94 64.20 69.37 68.23 Part Model 80.38 78.22 79.30 54.54 67.30 69.73 79.45 72.70
5.1 Compute vs Memory Reduction
A key aspect of SwiftKV is combining prefill compute reduction and KV cache compression (AcrossKV). While many prior works address KV cache compression alone, they are only effective when GPU memory is limited, and are less impactful on datacenter GPUs (e.g., A100 and H100) with sufficient memory and inference is compute-bound.
To illustrate, we construct an “ideal” KV compression scheme, where every layer’s KV cache is merged into a single layer (Merge-all-Layers). We retain the computation for all KV operations (i.e., $W_{kv}^{T}X$) but eliminate the memory for all layers > 1, leading to a single layer of KV cache. Merge-all-Layers represents a “best case compression scenario” with (1) extreme compression ratio beyond any published technique, e.g. 32×32$\times$32 × and 80×80$\times$80 × for Llama-3.1 8B-Instruct and 70B-Instruct, respectively, and (2) zero overhead, while most techniques (e.g., quantization, low-rank decomposition) add extra computations or data conversions.
Table 3 shows the throughput attained by Merge-all-Layers compared with the baseline model and its SwiftKV variants under various memory constraints. As shown, Merge-all-Layers outperforms only in very low memory scenarios (e.g. 16GB and 20GB) when there is barely enough memory for just the model weights, and is only marginally (10%) better than the baseline model when using all 80GB memory. On the other hand, SwiftKV attains 35% higher throughput than the baseline at 80GB even without AcrossKV. When combined with 4×4$\times$4 × AcrossKV using FP8-quantized KV cache, SwiftKV can approach the throughput of Merge-all-Layers even at a more limited 20GB of memory.
@@ 5.2 The Impact of Distillation
To demonstrate the effectiveness of our distillation method, we train Llama-3.1-8B-Instruct with 50% SwiftKV and no AcrossKV using the standard language model loss, and compare it with our distillation based approach discussed in  Sec. 3.4. The results are shown in Table 4 (a). As we can see, the model trained with distillation has a 2.64 point higher average. Particularly, for generative tasks, i.e., MMLU-Cot and GSM-8K, the performance improvement is 4.13 and 6.74, respectively.
@@
Our distillation method only fine-tuned the ${\mathbf{W}}_{QKV}$ parameters hypothesizing that this preserves the original model’s knowledge better than full model fine-tuning. This aligns with Meng et al. (2024), Geva et al. (2021), and Elhage et al. (2021), which suggest that MLP layers player a more prominent role in storing knowledge.
To validate this, we fine-tuned a model with 50% SwiftKV on Llama-3.1-8B-Instruct where all parameters in the latter 50% of layers are trained. The results are shown in Table 4 (b). The model quality of full model distillation is about 4.5 points lower than our proposed partial model distillation.
@@ 6 Conclusions
We presented SwiftKV, a model transformation for reducing inference cost for prefill-dominant workloads, combined with a KV cache reduction strategy to reduce memory footprint, and a light-weight distillation procedure to preserve model accuracy. SwiftKV demonstrates strong results and leaves room for exploration in parameter-preserving transformations to further optimize inference.
Appendix A Main Experiment Details
A.1 Training and Quality Evaluation Details
For datasets, we use a mixture of HuggingFaceH4/ultrachat_200k, teknium/OpenHermes-2.5, and Open-Orca/SlimOrca which totals around 680M tokens. We set training epochs to be 2, learning rate to be 3e-4, weight decay to be 0.05, warm up ratio to be 5%, maximum sequence length to be 8192 with attention separated sequence packing, the distillation temperature to be 2.0.
Our evaluation follows https://huggingface.co/neuralmagic/Meta-Llama-3.1-8B-Instruct-FP8 using the github repository https://github.com/neuralmagic/lm-evaluation-harness/tree/llama_3.1_instruct. The main reason behind this is that the implementation implemented chat-templated evaluations for several of our evaluation tasks, which is especially important for the Llama-3.1/3.2 models. For all tasks, we follow the same number of few shots and/or chain of thoughts as the provided commands. We present the number of shots and metrics used in the paper in Table A.1.
Table A.1: The setting for different tasks
@@ A.1 Training and Quality Evaluation Details
A.2 Inference Speedup Evaluation Details
@@
A.3 Llama-3.1-8B-Instruct Latency Results
See Fig. A.1.
Figure A.1: Time to first token (TTFT, top) and time per output token (TPOT, bottom) for input lengths 2000 (left), 8000 (middle), and 32000 (right) for Llama-3.1-8B-Instruct. For each experiment, a range of different request arrival rates is simulated. Each request generates 256 output tokens.
A.4 Inference Results with SGLang
In addition to vLLM, we also implemented SwiftKV on SGLang Zheng et al. (2024). SGLang differs from vLLM in that it leverages RadixAttention and Prefix Caching as first-class citizens, but otherwise supports many of the same features as vLLM, such as chunked-prefill Agrawal et al. (2024); Holmes et al. (2024).
We report the throughput results using SGLang in Table A.2. Overall, we observe similar relative improvements over the baseline (1.4 – 1.8$\times$ higher throughput for Llama-3.1-8B-Instruct, and 1.5 – 1.8$\times$ for Llama-3.1-70B-Instruct) using SGLang as vLLM (Fig. 3).
Table A.2: Inference throughput for Llama-3.1-8B-Instruct and Llama-3.1-8B-Instruct on SGLang.
Model Input length Output length Baseline (tokens/s) 50% SwiftKV (tokens/s) 50% SwiftKV + 4$\times$ AcrossKV (tokens/s) Llama-3.1-8B-Instruct 2000 256 27.4K 36.2K 38.9K 8000 256 22.9K 31.0K 34.0K 32000 256 16.9K 25.9K 26.6K 128000 256 7.66K 13.2K 14.0K Llama-3.1-70B-Instruct 2000 256 11.6K 15.7K 17.3K 8000 256 10.8K 16.1K 17.8K 32000 256 8.82K 14.0K 15.3K 128000 256 4.78K 8.21K 8.75K
A.5 Inference Results on ShareGPT
We provide additional evaluations using the ShareGPT dataset ShareGPT Team (2023), which consists of real-world conversations between users and ChatGPT OpenAI (2022). To better match our own observed request lengths (i.e. inputs $\geq$ 10$\times$ outputs), and to cover a broader range of scenarios, we also benchmark different versions of ShareGPT filtered by minimum input/output ratios. These datasets preserve the internal diversity of request lengths from ShareGPT. We report the average input/output length ratios and the measured performance for each of these filtered datasets below.
Table A.3 shows the results. Overall, we observe similar percentage improvements from SwiftKV as our main synthetic-dataset experiments, i.e. 1.25 – 1.7$\times$ and 1.25 – 1.8$\times$ higher throughput for Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct respectively for average length ratios up to $\approx$ 100 (similar ratio to the 32K input length experiments in Fig. 3).
Table A.3: Inference throughput for Llama-3.1-8B-Instruct and Llama-3.1-8B-Instruct on ShareGPT.
Model Min length Avg length ratio of Baseline 50% SwiftKV 50% SwiftKV + 4$\times$ AcrossKV ratio filter filtered dataset (tokens/s) (tokens/s) (tokens/s) Llama-3.1-8B-Instruct 0 (Original) 1.5 23.7K 27.6K 29.4K 0.2 3.4 25.8K 31.3K 31.9K 1 6.5 27.2K 35.1K 37.3K 2 10 30.3K 41.5K 43.7K 10 26 37.1K 54.7K 56.6K 20 40 37.7K 57.6K 59.9K 100 150 40.3K 64.2K 67.0K Llama-3.1-70B-Instruct 0 (Original) 1.5 9.73K 11.2K 12.2K 0.2 3.4 10.4K 13.2K 14.2K 1 6.5 11.4K 15.6K 16.0K 2 10 12.6K 18.0K 19.0K 10 26 14.1K 22.6K 23.2K 20 40 14.1K 22.9K 24.1K 100 150 14.6K 24.9K 25.8K
Appendix B Additional Ablations and Discussions
B.1 Combining KV Compression Methods
Table B.1: Llama-3.1-8B-Instruct KV cache quantization results.
SwiftKV operates in an orthogonal design space to other KV compression methods and can be combined with techniques such as sliding window (Jiang et al., 2023), token-level pruning (Liu et al., 2024c) and quantization (Hooper et al., 2024). We show the combined effect of SwiftKV with per-token KV cache FP8 quantization  (Yao et al., 2022).  Table B.1 shows the accuracy degradation is within 0.4 points for all cases, even though we applied post-training quantization with no quantization-aware finetuning.
B.2 Inter-layer AcrossKV vs Intra-Layer KV cache Reduction